OpenStudy (anonymous):

write the expression as a single logarithm. log3+log5

OpenStudy (luigi0210):

\[\LARGE loga+logb=\log(a*b)\]

OpenStudy (anonymous):

One of the log rules is: log (ab) = log a + log b log (3 times 5) = log3 + log5 log 15 would be your answer.

OpenStudy (luigi0210):

And \[\LARGE loga-logb=\log\frac{a}{b}\]

OpenStudy (anonymous):

what about 2 log7+ log 8

OpenStudy (anonymous):

is it 2 log56?

OpenStudy (luigi0210):

Another rule when condensing: \[\LARGE 3logz=logz^3\]

OpenStudy (luigi0210):

The factor at the end returns and becomes an exponent, and vice versa

OpenStudy (anonymous):

so the answer is logz^3?

OpenStudy (luigi0210):

What? no..

OpenStudy (anonymous):

what is it?

OpenStudy (luigi0210):

I'm not here to give answers..

OpenStudy (anonymous):

Rileymarie6666, we are supposed to explain how to do the question so that you can understand how to do it yourself. We're not supposed to just give you answers. That wouldn't really help you learn. There are log rules you need know so that you can do these questions. 2 log(7) + log(8) is your question. You can change 2log(7) into log (7^2) because of the law rule that Luigi0210 explained. Now you have log(49) + log(8). Do you remember the rule you used to do the first question?